What is the measure of a base angle of an isosceles triangle if the vertex angle measures 38° and each of the two congruent sides measures 21 units?
Answer by jim_thompson5910(35256)

Each base angle is (180-38)/2 = 142/2 = 71 degrees

Note: the info "each of the two congruent sides measures 21 units" is not needed in this case.